OpenStudy (anonymous):

Lee started making contributions to a Traditional IRA when he got his first job, at age 28. His contributions averaged $2,900 annually. Lee was in the 30% tax bracket during his working years, when he continued to make these annual contributions. The average annual rate of return on the account was 5.7%. Upon retirement at age 71, Lee stopped working and making IRA contributions. Instead, he started living on withdrawals from the retirement account. At this point, Lee dropped into the 20% tax bracket. Factoring in taxes, what is the effective value of Lee’s Traditional IRA at retirement? Assum

OpenStudy (anonymous):

7. Lee started making contributions to a Traditional IRA when he got his first job, at age 28. His contributions averaged $2,900 annually. Lee was in the 30% tax bracket during his working years, when he continued to make these annual contributions. The average annual rate of return on the account was 5.7%. Upon retirement at age 71, Lee stopped working and making IRA contributions. Instead, he started living on withdrawals from the retirement account. At this point, Lee dropped into the 20% tax bracket. Factoring in taxes, what is the effective value of Lee’s Traditional IRA at retirement? Assume annual compounding. (1 point)$441,405.34 $400,703.58 $151,480.27 $350,615.64

ganeshie8 (ganeshie8):

use the same formula and find the accumulated sum first

ganeshie8 (ganeshie8):

\(\large FV = C\bullet \dfrac{(1+i)^{nt}-1}{i}\)

OpenStudy (anonymous):

hmm one sec. ill plug it in

ganeshie8 (ganeshie8):

Once you have the accumulated sum, simply subtract the taxes...

ganeshie8 (ganeshie8):

take ur time :)

OpenStudy (anonymous):

what did i do wrong ? :/ 2900*( ((1+0.057)^(43)-1)/(0.057) ) - 43*.30*2900

ganeshie8 (ganeshie8):

Very good try !

OpenStudy (anonymous):

-_- math sucks as of today

ganeshie8 (ganeshie8):

In traditional IRA, you need to cut the taxes on accumulated sum.

ganeshie8 (ganeshie8):

accumulated sum = 2900*( ((1+0.057)^(43)-1)/(0.057) ) right ?

OpenStudy (anonymous):

wat do you mean on the second part ?

ganeshie8 (ganeshie8):

when he was retired, he fell into 20% tax bracket. so you need to cut 20% taxes on this sum : effective value of money = ( 2900*( ((1+0.057)^(43)-1)/(0.057) ) ) * (1- 0.2)

ganeshie8 (ganeshie8):

after cutting that 20% tax on accumulated sum, u wud get : http://www.wolframalpha.com/input/?i=%282900*%28+%28%281%2B0.057%29%5E%2871-28%29-1%29%2F%280.057%29+%29+%29*%281-.2%29

OpenStudy (anonymous):

dang. Lol. Hmmmmmmmmmmmm ONNNNNNNNNNNNNNNNNNNNE Final one ? :P

ganeshie8 (ganeshie8):

remember this : Traditional IRA : taxed on accumulated sum AFTER retirement Roth IRA : taxed when u work itself when u make payments BEFORE retirement

ganeshie8 (ganeshie8):

sure :)